OpenStudy (anonymous):

What is the equation of the line that passes through the points (−2, 1) and (1, 10)?

OpenStudy (anonymous):

In point slope, the equation is\[(y-1) = 3(x+2)\]

OpenStudy (anonymous):

Right I understand how do you do it though

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

first you start off by determining the slope of the line then you can put it in point slope form and convert it into slope intercept form or standard form depending on what hte question is asking for personally i usually write line equations in slope intercept form as the final form

OpenStudy (anonymous):

ok but how do you determine the slope

OpenStudy (anonymous):

The general formula for finding slope (m) is in his post, but here it is anyway: \[\frac{ y_2 - y_1 }{ x_2 - x_1 }\] Or, as I like to say, rise over run (but that's kinda confusing and elementary)

OpenStudy (anonymous):

ok so the slope would be 3/9 or 1/3

OpenStudy (anonymous):

The slope would be 3. \[\frac{ 10-1 }{ 1+2 } = \frac{ 9 }{ 3 } = 3\]

OpenStudy (anonymous):

right i flipped it

OpenStudy (anonymous):

Can anybody help me

OpenStudy (anonymous):

A solution calculated with Mathematica 9 is attached.

OpenStudy (anonymous):

Thanks so muck robotobey